Welcome >  Stations >  Herstmonceux, United Kingdom (7840) >  Station Logs

Station History Log - Herstmonceux, United Kingdom (7840)

SOD Year Day
of Year
Time
of Day
Data Impact System Description
78403501 2022 041 11:00 1 9.01.02 H-Maser providing timing signals again following repair and testing
78403501 2021 322 14:00 1 5.01 kHz laser returns to operation
78403501 2021 298 16:00 1 5.01 kHz laser problem, switch to 12Hz laser.
78403501 2021 146 11:00 1 9.01.02 H-Maser problem, timing signals switched to S650 source
78403501 2020 269 10:00 1 5.01 Switched to kHz laser for SLR operations
78403501 2020 240 09:00 0 5.01 12Hz laser service
78403501 2020 211 19:00 1 5.01 Switched to older 12Hz laser for SLR operations
78403501 2020 198 18:00 1 06.01.07 New SPAD cable fitted
78403501 2020 049 18:00 0 7.02 New Ranging PC installation complete
78403501 2020 041 21:00 0 7.02 PC fatal crash. Motherboard stopped working.
78403501 2019 136 08:30 1 05.02 New kHz freq-doubler crystal fitted
78403501 2019 098 08:30 1 05.02 kHz chiller water changed and flow cleaned
78403501 2018 339 08:30 1 9.01.02 GPS reference replaced with new S650 GPS receiver
78403501 2018 275 08:30 1 06.01.07 New SPAD cable fitted
78403501 2018 268 08:30 0 99 correction to rms in CRD record 50 pass statistics
78403501 2018 254 08:30 0 04.01 Narrow spectral filter fitted with oven. Tilted
78403501 2018 141 08:30 1 05 Coude mirrors cleaned
78403501 2018 120 08:30 0 04.01 New narrow spectral filter fitted
78403501 2018 116 08:30 0 99 Mains electrical provided by power grid
78403501 2018 099 08:30 0 99 Mains electrical provided by generator
78403501 2018 074 08:30 0 04.01 New daytime narrow filter assemby fitted
78403501 2018 036 08:30 1 08.02 New external target adopted as primary calibration. New survey values used.
78403501 2018 002 08:30 1 9.01.02 Resynced H-Maser with GPS3 1pps
78403501 2017 362 08:30 1 9.01.02 Fault with H-Maser detected and reset
78403501 2017 328 08:30 1 05.02 kHz start diode adjusted to give 2V ouput
78403501 2017 327 08:30 0 05.02 kHz laser service
78403501 2017 319 08:30 1 9.01.02 Fault with H-Maser detected, previous 10 days SLR data corrected and resubmitted.
78403501 2017 311 08:30 1 9.01.02 PC clock and ISA clock drift detected
78403501 2017 160 08:30 1 9.01.02 PC clock driven by 5MHz maser signal
78403501 2017 097 15:30 1 5.01 Switched to kHz laser for SLR operations
78403501 2017 090 08:30 1 5.01 Switched to older 12Hz laser for SLR operations
78403501 2017 090 08:30 1 5.03 Water flow rate through kHz laser found to be very low and not providing sufficient cooling
78403501 2017 083 08:30 1 9.01.02 SLR timing switched back to H-Maser for 10MHz and 1pps
78403501 2017 071 08:30 1 9.01.02 H-Maser failure. SLR timing switched to GPS3 for 10MHz and 1pps
78403501 2016 244 08:30 0 10. Reduction in return rate bin size to identify high signal strength data for removal
78403501 2016 239 08:30 0 6.01.07 Adjustement to real-time ND filter control to reduce signal strength
78403501 2016 141 08:30 1 9.01.02 1pps cable to HxET replaced
78403501 2016 076 08:30 0 4.01 Spare daytime narrow filter installed
78403501 2015 279 08:30 0 04.01 Previous daytime narrow filter setup reinstalled
78403501 2015 247 08:30 0 05.01 1/4 wave plate installed for kHz laser to provide circular polarisation (this was unsuccessful)
78403501 2015 237 08:30 0 04.01 New daytime narrow filter setup, including new narrow oven filter, installed
78403501 2015 189 08:30 0 04.01 New daytime narrow filter setup, including new blocking filter, installed
78403501 2015 145 08:30 1 06.01.07 Start and Stop pulses to Riga and HxET timers reduced in duration to 30ns
78403501 2014 279 08:30 1 05.02 kHz laser upgrade to Nd:YAG with increased pulse energy and 1kHz ability
78403501 2014 056 08:30 0 99 Electrical rewiring in control room, offices and in telescope dome.
78403501 2014 035 08:30 2 05.01 Nd:YAG start diode and discriminator adjusted to improve calibration leading edge.
78403501 2014 027 08:30 1 05.01 Nd:YAG laser service and pulse selector trial. Completed 3/2/14 and selector removed.
78403501 2013 221 08:30 1 05 Emitter end optic cleaned
78403501 2013 211 08:30 0 05 Coude mirrors M2, M3 and M4 replaced
78403501 2013 207 08:30 1 08.01 TR reduction software adjusted for better performance with kHz signal to noise
78403501 2013 137 08:30 1 05.02 2kHz laser start diode adjustment
78403501 2013 130 08:30 1 06.01.07 Gate cable to SPAD changed
78403501 2013 128 08:30 1 05.02 2kHz Nd:VAN laser back in operation
78403501 2013 008 08:30 2 08.01 New power supply fitted to Lecroy discriminators to eliminate cause of calibration jumps
78403501 2012 289 08:30 1 08.01 SPAD signals in Lecroy discriminator changed
78403501 2012 282 08:30 1 08.01 Change in Lecroy discriminator setup
78403501 2011 280 08:30 1 06.01.07 Cables to event timer changed
78403501 2011 264 08:30 1 05.01 Nd:YAG laser service
78403501 2011 255 08:30 1 06.01.07 SPAD5 removed and C-SPAD fitted
78403501 2011 255 08:30 1 06.01.07 Cables to event timer swapped
78403501 2011 252 08:30 1 06.01.07 SPAD signal cable changed
78403501 2011 244 08:30 1 06.01.07 SPAD signal and gate cables switched
78403501 2011 110 08:30 0 06 Spectral filter switch and cable replaced
78403501 2011 053 08:30 1 05.01 Nd:YAG laser service
78403501 2011 033 08:30 1 06.01.07 SPAD signal and gate cables changed
78403501 2010 302 08:30 1 05 Coude mirrors M2 swapped for old M1
78403501 2010 279 08:30 1 99 PSU damaged by powercut and replaced
78403501 2010 250 08:30 2 12.02.03 New Rotronics Met device becomes primary reference
78403501 2010 245 08:30 1 06 New Telescope dichroic mirror fitted
78403501 2010 145 08:30 1 05.01 Nd:YAG laser service
78403501 2010 111 08:30 1 06 Old Telescope dichroic mirror refitted
78403501 2010 081 08:30 0 05.02 2kHz laser maintenance
78403501 2010 049 08:30 1 06.01.07 SPAD5 fitted
78403501 2010 036 08:30 1 05.02 2kHz laser start diode adjustment
78403501 2010 012 08:30 1 06 Telescope dichroic mirror removed for glass replacement
78403501 2010 004 08:30 1 06.01.07 SPAD5 removed and C-SPAD installed
78403501 2009 280 08:30 1 05 Emitter beam expander exit window replaced
78403501 2009 168 08:30 1 05.01 Nd:YAG start diode removed and refitted
78403501 2009 096 17:00 1 06.01.07 New SPAD cables fitted
78403501 2009 007 08:30 1 06.01.07 SPAD5 fitted
78403501 2008 353 08:30 1 06.01.07 SPAD5 tests
78403501 2008 333 08:30 1 06.01.05 SR620 reinstalled in system
78403501 2008 331 08:30 1 05.02 2kHz Nd:VAN operational, new freq-doubler crystal fitted
78403501 2008 308 08:30 1 05.02 2kHz Nd:VAN offline
78403501 2008 294 08:30 0 05 Coude mirror M3 replaced
78403501 2008 290 08:30 1 06.01.07 SPAD signal cable changed
78403501 2008 280 08:30 0 05.02 2kHz Nd:VAN maintenance
78403501 2008 246 08:30 0 05.01 Nd:YAG maintenance
78403501 2008 189 08:30 1 06.01.07 C-SPAD fitted
78403501 2008 177 08:30 2 13.01 IGN site-survey
78403501 2008 144 08:30 0 05.01 Nd:YAG maintenance
78403501 2008 133 08:30 1 05.01 Nd:YAG interpulse gap changed
78403501 2008 121 08:30 1 06.01.07 C-SPAD damaged. Replaced with SPAD-3B
78403501 2008 055 08:30 0 05 Coude mirror M1 replaced
78403501 2008 029 08:30 1 05.01 Nd:YAG laser part service
78403501 2007 347 08:30 0 05.02 2kHz Nd:VAN maintenance
78403501 2007 307 08:30 2 08.01 DG535 left in wrong state after tests, corrected after 1 day.
78403501 2007 284 08:30 0 05.02 2kHz Nd:VAN start diode moved
78403501 2007 281 08:30 0 05.02 2kHz Nd:VAN maintenance
78403501 2007 250 08:30 1 06.01.07 C-SPAD fitted, output sent through Lecroy
78403501 2007 229 08:30 0 08.01 Cable change, Lecroy taken out of stop chain
78403501 2007 227 08:30 0 08.01 Adjustment to cables used
78403501 2007 218 08:30 1 08.01 Start signal moved to different Lecroy discriminator block
78403501 2007 204 08:30 1 05.01 New Nd:YAG start diode and cable fitted
78403501 2007 198 08:30 0 05.01 Nd:YAG maintenance
78403501 2007 187 08:30 1 06.01.07 SPAD-3B installed for tests
78403501 2007 102 08:30 0 05.01 Nd:YAG maintenance
78403501 2007 022 08:30 0 05.01 Nd:YAG maintenance

Find more topics on the central web site of the Technical University of Munich: www.tum.de